# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ IV-C Algorithm
We are now ready to outline the dual subgradient method with the proposed initialization and the stepsize rule that leverages FGOR characteristics. In the rest of this section, either the ray $\mathcal{R}^{\star}(\bar{\mathbf{y}})$ [cf.  (25)] or the line $\mathcal{L}^{\star}(\bar{\mathbf{y}})$ [cf.  (27)] is considered to be available. However, the algorithm is outlined with $\mathcal{L}^{\star}(\bar{\mathbf{y}})$ for clarity  [^1] If $\mathcal{R}^{\star}(\bar{\mathbf{y}})$ is considered, the algorithm remains the same, except that $\mathcal{L}^{\star}$ is replaced by $\mathcal{R}^{\star}$., cf.  Algorithm 1.
Algorithm 1 Dual Subgradient Method
1:$\Lambda(\bar{\mathbf{y}})={\mathcal{L}}^{\star}(\bar{\mathbf{y}})$.
2:$\bm{\lambda}^{(0)}\in\Lambda(\bar{\mathbf{y}})$.
3:$k=0$.
4:if $\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}^{(k)}{\in}\partial f_{0}(\bar{\mathbf{y}})+\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}})$ then $\triangleright$ in FGOR region
5:     Perform (14) with $\gamma_{k}=\gamma$ and $\mathbf{s}^{(k)}=-\bm{\mu}$.
6:     Set $k:=k+1$ and go to step 2.
7$\triangleright$ FGOR region is exited
8:     Go to step 8. $\triangleright$ to an existing stepsize rule
9:end if
10:repeat
11:     Perform (14) with an existing stepsize rule.
12:     $k:=k+1$.
13:until a stopping criterion true
The condition to be checked at step 2 of Algorithm 1 is unfavorable in a distributed optimization setting. Therefore, we may wish to have at our disposal an alternative, yet equivalent representation that is favorable in a distributed setting. This is the subject of the following remark.
@@
$\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}^{(k)}\in\partial f_{0}(\bar{\mathbf{y}})+\partial\delta_{\mathcal{Y}}(\bar{\mathbf{y}})\iff\mathbf{y}^{(k)}=\bar{\mathbf{y}}$ (28)
@@
In the sequel, we place a greater emphasis on the application of FGOR characteristics to the global consensus problem [49, § 7] since it is a widely used problem formulation in numerous application domains [50, 6, 4, 51, 52].
@@ V Global Consensus Problem
In this section, we explore how FGOR characteristics [cf.  Proposition 1, Corollary 1, § IV] can be exploited when solving the global consensus problem. In particular, we consider the problem
$\begin{array}[]{ll}\mbox{minimize}&\sum_{i{=}1}^{m}f_{i}(\mathbf{z})\\ \mbox{subject to}&\mathbf{z}\in\mathcal{Z},\end{array}$ (29)
where the variable is $\mathbf{z}\in{\rm I\!R}^{n}$. We assume that the functions $f_{i}:{\rm I\!R}^{n}\rightarrow{\rm I\!R}$, $i=1,\ldots,m$ and the constraint set $\mathcal{Z}\subseteq{\rm I\!R}^{n}$ are conforming to Assumption 1 and Assumption 2. A commonly used distributed solution method for solving (29) is based on dual decomposition [40].
@@ V-A The Standard Dual Decomposition Method
Problem (29) can be equivalently reformulated as follows:
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})\\ \mbox{subject to}&\mathbf{y}_{i}\in\mathcal{Z},\ i\in\mathcal{S}\\ &\mathbf{y}_{i}=\mathbf{y}_{i+1},i\in\mathcal{S}\setminus\{m\},\end{array}$ (30)
@@
Algorithm 2 Dual Decomposition Algorithm for Global Consensus Problem
$\bm{\lambda}\in{\rm I\!R}^{n(m-1)}$.
@@
7:     CS computes ${\mathbf{s}}^{(k)}{=}[(\mathbf{y}^{(k)}_{1}{-}{\mathbf{y}}^{(k)}_{2})^{\mbox{\scriptsize T}}{\ldots}(\mathbf{y}^{(k)}_{m{-}1}{-}\mathbf{y}^{(k)}_{m})^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$.
8:     CS computes $\bm{\lambda}^{(k+1)}=\bm{\lambda}^{(k)}+\gamma_{k}\mathbf{s}^{(k)}$.
@@
In Algorithm 2, CS performs the dual variable update [cf.  step 7] using any standard stepsize rule [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. It is worth noting that when the FGOR characteristic of $g$ is utilized, the communication steps 5 and 8 in Algorithm 2 are not needed as long as $\bm{\lambda}^{(k)}$ resides in a FGOR region. In particular, utilizing the FGOR characteristic of $g$, we can improve the performance of Algorithm 2 not only in terms of the speed of the convergence but also in terms of communication efficiency. We discuss this idea more formally in the following subsection.
V-B Using the FGOR Characteristic of $g$ on Algorithm 2
@@
We let $\mathcal{Z}^{m}=\mathcal{Y}$ for clarity. Suppose that the feasibility problem (26) associated to problem (30) is feasible for some $\mathbf{\bar{y}}=[\mathbf{\bar{y}}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \mathbf{\bar{y}}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}\in\texttt{bnd}~{}\mathcal{Y}$, where $\mathbf{\bar{y}}_{i}\in\mathcal{Z}$, $\forall i\in\mathcal{S}$. Then Algorithm 3 outlines how the FGOR characteristic of the dual function $g$ is leveraged to solve problem (30) using dual decomposition.
Algorithm 3 Dual Decomposition Algorithm with FGOR
1:$\Lambda(\bar{\mathbf{y}})={\mathcal{L}}^{\star}(\bar{\mathbf{y}})$.
2:$\bm{\lambda}^{(0)}\in\Lambda(\bar{\mathbf{y}})$.
3:$k=0$.
4:$\forall i$, agent $i$ computes:
@@
6:CS computes $\bar{b}=\max_{i\in\mathcal{S}}b_{i}$.
7:$\forall i$, CS broadcasts $\bar{b}$ to agent $i$.
8:if $\bar{b}=0$then $\triangleright$ in FGOR region
9:     $\forall i$, agent $i$ and CS perform (14) with
$\gamma_{k}=\gamma\ \mathrm{and}\ \mathbf{s}^{(k)}=-\bm{\tilde{\mu}}.$
10:     Set $k:=k+1$ and go to step 2.
11$\triangleright$ FGOR region is exited
12:     Go to step 12.
13:end if
14:$\forall i$, agent $i$ transmits $\mathbf{y}_{i}^{(k)}$ to CS.
15:CS performs (14) using $\gamma_{k}$ in Algorithm 2, where
${\mathbf{s}}^{(k)}{=}\big{[}\big{(}\mathbf{y}^{(k)}_{1}{-}{\mathbf{y}}^{(k)}_{2}\big{)}^{\mbox{\scriptsize T}}\ldots\big{(}\mathbf{y}^{(k)}_{m{-}1}{-}\mathbf{y}^{(k)}_{m}\big{)}^{\mbox{\scriptsize T}}\big{]}^{\mbox{\scriptsize T}}.$
16:Switch to Algorithm 2 with $\bm{\lambda}=\bm{\lambda}^{(k+1)}$.
Note that Algorithm 3 is a blend of Algorithm 1 [cf.  § IV] and Algorithm 2 in the sense that we exploit the steps described in Algorithm 1 to leverage FGOR characteristics on Algorithm 2. In particular, steps 2-6 of Algorithm 3 are corresponding to step 2 of Algorithm 1, where step 6 is the criterion we set in Algorithm 3 to check whether $\bm{\lambda}^{(k)}$ is in a FGOR region. Roughly speaking, if $\bar{b}=0$, it is straightforward to see that $\mathbf{y}^{(k)}=\mathbf{\bar{y}}$ [cf.  (34)], and thus $\bm{\lambda}^{(k)}$ is in a FGOR region, cf.  Remark 4. Moreover, steps 3-4 and steps 5-11 of Algorithm 1 are corresponding to steps 7-8 and steps 9-14 of Algorithm 3, respectively.
A natural selection for $\bm{\lambda}^{(0)}$ in Algorithm 3 is $\bm{\lambda}^{(0)}=\bm{\tilde{\mu}}$, cf.  27. It is worth noting that each agent $i$ and the CS can individually solve (26). If feasible, the solutions $\bm{\tilde{\eta}}$ and $\bm{\tilde{\mu}}$ obtained by solving (26) are the same for all agents and CS. Thus, both CS and each agent $i$ know $\bm{\lambda}^{(0)}$ and the dual subgradient is $\mathbf{s}^{(k)}{=}-\bm{\tilde{\mu}}$ as long as $\bm{\lambda}^{(k)}$ resides in the respective FGOR region, cf.  steps 6-7. This enables each agent $i$ to perform the dual variable update (14) individually and in parallel until $\bm{\lambda}^{(k)}$ exits the FGOR region, cf.  step 9. However, the CS also performs the same computation to keep track of the update $\bm{\lambda}^{(k)}$ to use it after the algorithm steps exit the FGOR region, cf.  steps 13-14. We discuss the efficiency of Algorithm 3 compared to Algorithm 2 in the sequel in detail.
@@ V-B1 Fast Convergence with FGOR
Rather than choosing $\bm{\lambda}^{(0)}$ arbitrarily as in Algorithm 2, Algorithm 3 is initialized with $\bm{\lambda}^{(0)}$ by exploiting the FGOR characteristic of the dual function $g$, cf.  problem (26). Unlike Algorithm 2, $\bm{\lambda}^{(0)}$ in Algorithm 3 permits one to use fixed stepsizes and makes the subgradient algorithm [cf.  (14)] evolve along the specified ray yielding constant increments of $g$ as long as $\bm{\lambda}^{(k)}$ resides in a FGOR region. Intuitively speaking, such a choice is fast in the sense that it directly drives $\bm{\lambda}^{(k)}$ toward $\bm{\lambda}^{\star}$ instead of potential detours where diminishing stepsizes such as $\gamma_{0}/k$, $\gamma_{0}/\sqrt{k}$ can be ineffective. Our choice of $\bm{\lambda}^{(0)}$ for Algorithm 3 can also be considered as a warm-start initialization that expedites the convergence of the algorithm toward the intended solution.
@@ V-B2 Efficient Communication with FGOR
Note that Algorithm 2 requires each agent $i$ to transmit $n$ real numbers, i.e., $\mathbf{y}_{i}^{(k)}\in{\rm I\!R}^{n}$ in step 5. In addition, it also requires CS to broadcast $n(m-1)$ real numbers to agents, cf.  step 8. Therefore, if $b$ bits are used to represent each real number, then Algorithm 2 requires $mnb+n(m-1)b=nb(2m-1)$ bits per epoch  [^2] In the case of $64$-bit double precision floating point number format $b=64$.. However, at the corresponding steps of Algorithm 3 [cf.  step 3 and step 5], only one-bit communication is needed and it does not depend on the length $n$ of $\mathbf{y}_{i}^{(k)}$. More specifically, in each epoch, Algorithm 3 requires only $m+1$ bits, i.e., $m$ bits at step 3 and one bit at step 5, as long as $\bm{\lambda}^{(k)}$ resides in a region of FGOR. It is worth noting that engineering problems in real-world applications can consist of decision variables whose size $n$ can be in the order of several thousands, or even more [53]. Thus, by leveraging FGOR characteristics, Algorithm 3 can yield a considerable reduction in the communication overhead.

### v3

@@ IV-C Algorithm
We solve (28) and choose the initialization point $\bm{\lambda}^{(0)}$ for the dual subgradient method (17) as follows:
$\bm{\lambda}^{(0)}=\hat{\bm{\lambda}}=c\bm{\tilde{\mu}}$, (30)
where $c\in{\rm I\!R}$, cf.  (29). Then the corresponding algorithm, together with the proposed stepsize rule [cf.  (19), (20)], is outlined in Algorithm 1. Note that the proposed $\bm{\lambda}^{(0)}$ [cf.  (30)] serves as a warm start for the dual subgradient method (17), as it enables (17) to use a constant stepsize and produce constant increments in the dual function $g$ as long as $\bm{\lambda}^{(k)}\in\mathcal{R}^{\star}(\mathbf{\bar{y}})$ [cf.  steps 2-3] allowing $\bm{\lambda}^{(k)}$ to quickly approach the dual optimal solution $\bm{\lambda}^{\star}$.
Algorithm 1 Dual Subgradient Method with FGOR
1:the solution $\tilde{\bm{\mu}}$ of problem (28), $c>0$, and $\beta>0$.
[…]
3:if $\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}^{(k)}{\in}\partial f_{0}(\bar{\mathbf{y}})+N_{\mathcal{Y}}(\bar{\mathbf{y}})$ then $\triangleright$ in RFGOR
4:  Perform (17) with $\gamma_{k}=\gamma$ and $\mathbf{s}^{(k)}=\beta\bm{\tilde{\mu}}$.
5:  Set $k:=k+1$ and go to step 2.
[…]
7:  Go to step 8. $\triangleright$ to an existing stepsize rule
8:end if
9:repeat
10:  Perform (17) with an existing stepsize rule.
11:  $k:=k+1$.
12:until a stopping criterion true
Note that the condition to be checked at step 2 of Algorithm 1 is unfavorable in a distributed optimization setting. Therefore, we may wish to have at our disposal an alternative, yet equivalent representation that is favorable in a distributed setting. This is the subject of the following remark.
@@
$\mathbf{A}^{\mbox{\scriptsize T}}\bm{\lambda}^{(k)}\in\partial f_{0}(\bar{\mathbf{y}})+N_{\mathcal{Y}}(\bar{\mathbf{y}})\iff\mathbf{y}^{(k)}=\bar{\mathbf{y}}$ (31)
@@
We establish next the convergence of Algorithm 1.
Remark 5
Suppose Assumption 1, Assumption 2, and Assumption 3 hold. Then, Algorithm 1 converges to an optimal solution $\bm{\lambda}^{\star}$ of the dual problem (16).
Proof:
If $\nabla g(\bm{\lambda}^{(0)})=\mathbf{0}$, then $\bm{\lambda}^{(0)}$ is an optimal solution of problem (16), and hence the convergence is trivial. Otherwise, either 1) $\bm{\lambda}^{(0)}\notin\textrm{RFGOR}$ [cf.  step 5] or 2) $\bm{\lambda}^{(0)}\in\textrm{RFGOR}$ [cf.  step 2]. In the first case, the control goes to step 6, after which the algorithm performs the standard subgradient method (17) with an existing stepsize rule [cf.  steps 8-11], guaranteeing the convergence. In the latter case, Algorithm 1 performs the subgradient method (17) with a constant stepsize $\gamma$ and with the gradient $\beta\bm{\tilde{\mu}}$. In this respect, we note that there is a support to $g$ at all $\bm{\lambda}^{(k)}\in\textrm{RFGOR}$, denoted $h$, which is given by the affine function $h(\bm{\lambda})=-\beta\bm{\tilde{\mu}}^{\mbox{\scriptsize T}}\bm{\lambda}+c$ for some $c\in{\rm I\!R}$, which doesn’t depend on $\bm{\lambda}^{(k)}$. Thus, at each $\bm{\lambda}^{(k)}\in\textrm{RFGOR}$, $g(\bm{\lambda}^{(k)})=h(\bm{\lambda}^{(k)})$. Moreover, since $g$ is concave and has an optimal solution $\bm{\lambda}^{\star}$  [^1] According to the assumptions, primal problem (1) has a unique solution $\mathbf{y}$$^{\star}$ and strong duality holds., $h$ is an upper-bound on $g$ such that $h(\bm{\lambda})>g(\bm{\lambda})$ for some $\bm{\lambda}$. Thus, there must be an index $\bar{k}>0$ such that $h(\bm{\lambda}^{(k)})=g(\bm{\lambda}^{(k)})$ for all $k\leq\bar{k}$ and $h(\bm{\lambda}^{(\bar{k}+1)})>g(\bm{\lambda}^{(\bar{k}+1)})$. This implies that $\bm{\lambda}^{(\bar{k}+1)}\not\in\textrm{RFGOR}$, and thus the algorithm switches to the standard subgradient method, guaranteeing the convergence according to steps 5,6,8, and 9. ∎
In the sequel, we place greater emphasis on the application of FGOR to the global consensus problem [Boyd-Parikh-Chu-Peleato-Eckstein-2010, § 7] since it is a widely used formulation in numerous application domains [Chen_consensus_2012, Carlo-Machine-Learning, Yang-survey-of-distributed-optimization-2019, Ball_consensus_2023, halsted-Multiple-robots-2021].
@@ V Global Consensus Problem
Let us now consider the problem
$\begin{array}[]{ll}\mbox{minimize}&(1/m)\sum_{i{=}1}^{m}f_{i}(\mathbf{z})\\ \mbox{subject to}&\mathbf{z}\in\mathcal{Z},\end{array}$ (32)
where the variable is $\mathbf{z}\in{\rm I\!R}^{n}$. We assume that the functions $f_{i}:{\rm I\!R}^{n}\rightarrow{\rm I\!R}$, $i=1,\ldots,m$ and the constraint set $\mathcal{Z}\subseteq{\rm I\!R}^{n}$ are conforming to Assumption 1 and Assumption 2. A commonly used distributed solution method for solving (32) is based on dual decomposition [Boyd-EE364b-PrimDualDecomp-07].
@@ V-A The Standard Dual Decomposition Method
Problem (32) is equivalent to the global consensus problem
$\begin{array}[]{ll}\mbox{minimize}&f_{0}(\mathbf{y})=(1/m)\sum_{i{=}1}^{m}f_{i}(\mathbf{y}_{i})\\ \mbox{subject to}&\mathbf{y}_{i}\in\mathcal{Z},\ i\in\mathcal{S}\\ &\mathbf{y}_{i}=\mathbf{y}_{i+1},i\in\mathcal{S}\setminus\{m\},\end{array}$ (33)
@@
Algorithm 2 Dual Decomposition for Global Consensus
[…]
@@
7:  CS computes ${\mathbf{s}}^{(k)}{=}[(\mathbf{y}^{(k)}_{2}{-}{\mathbf{y}}^{(k)}_{1})^{\mbox{\scriptsize T}}{\ldots}(\mathbf{y}^{(k)}_{m}{-}\mathbf{y}^{(k)}_{m{-}1})^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}$.
8:  CS computes $\bm{\lambda}^{(k+1)}=\bm{\lambda}^{(k)}-\gamma_{k}\mathbf{s}^{(k)}$.
@@
In Algorithm 2, CS performs the dual variable update [cf.  step 7] using any standard stepsize rule [Polyak_Intro_Opt, goldstein_1962, armijo_1966, beck_2009, bello_2016, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Damek_2018, zhu_2019, aybat_2019, wang_2021, chen_2019, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020]. It is worth noting that when the FGOR characteristic of $g$ is utilized, the communication steps 5 and 8 in Algorithm 2 are not needed as long as $\bm{\lambda}^{(k)}$ resides in an RFGOR region. In particular, utilizing the FGOR characteristic of $g$, we can improve the performance of Algorithm 2 not only in terms of the speed of the convergence but also in terms of communication efficiency. We discuss this idea more formally in the following subsection.
V-B Using FGOR Characteristic of $g$ on Algorithm 2
@@
We let $\mathcal{Z}^{m}=\mathcal{Y}$ for clarity. Suppose that problem (28) associated to problem (33) is feasible for some $\mathbf{\bar{y}}=[\mathbf{\bar{y}}_{1}^{\mbox{\scriptsize T}}\ \ldots\ \mathbf{\bar{y}}_{m}^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}\in\texttt{bnd}~\mathcal{Y}$, where $\mathbf{\bar{y}}_{i}\in\mathcal{Z}$, $\forall i\in\mathcal{S}$  [^2] The feasibility of problem (28) associated to problem (33) is shown in § V-C, cf.  Lemma 2.. Then Algorithm 3 outlines how the FGOR characteristic of the dual function $g$ is leveraged to solve problem (33) using dual decomposition.
Algorithm 3 Dual Decomposition with FGOR for Global Consensus
1:the solution $\tilde{\bm{\mu}}$ of problem (28), $c>0$, and $\beta>0$.
[…]
@@
5:CS computes $\bar{b}=\max_{i\in\mathcal{S}}b_{i}$.
[…]
7:if $\bar{b}=0$, then $\triangleright$ in RFGOR region
8:  $\forall i$, agent $i$ and CS perform (17) with
$\gamma_{k}=\gamma\ \mathrm{and}\ \mathbf{s}^{(k)}=\beta\bm{\tilde{\mu}}$.
9:  Set $k:=k+1$ and go to step 2.
[…]
11:  Go to step 12.
12:end if
[…]
14:CS performs (17) using $\gamma_{k}$ in Algorithm 2, where
${\mathbf{s}}^{(k)}{=}[(\mathbf{y}^{(k)}_{2}{-}{\mathbf{y}}^{(k)}_{1})^{\mbox{\scriptsize T}}\ldots(\mathbf{y}^{(k)}_{m}{-}\mathbf{y}^{(k)}_{m-1})^{\mbox{\scriptsize T}}]^{\mbox{\scriptsize T}}.$
15:Switch to Algorithm 2 with $\bm{\lambda}=\bm{\lambda}^{(k+1)}$.
Note that Algorithm 3 is a blend of Algorithm 1 [cf.  § IV] and Algorithm 2 in the sense that we exploit the steps described in Algorithm 1 to leverage FGOR characteristics on Algorithm 2. In particular, steps 2-6 of Algorithm 3 are corresponding to step 2 of Algorithm 1, where step 6 is the criterion we set in Algorithm 3 to check whether $\bm{\lambda}^{(k)}$ is in an RFGOR region. Roughly speaking, if $\bar{b}=0$, it is straightforward to see that $\mathbf{y}^{(k)}=\mathbf{\bar{y}}$ [cf.  (37)], and thus $\bm{\lambda}^{(k)}$ is in an RFGOR region, cf.  Remark 4. Moreover, steps 3-4 and steps 5-11 of Algorithm 1 are corresponding to steps 7-8 and steps 9-14 of Algorithm 3, respectively.
[…]
@@ V-B1 Fast Convergence with FGOR
Rather than choosing $\bm{\lambda}^{(0)}$ arbitrarily as in Algorithm 2, Algorithm 3 is initialized with $\bm{\lambda}^{(0)}$ [cf.  (30)] by exploiting the FGOR characteristic of the dual function $g$, cf.  problem (28). Unlike Algorithm 2, $\bm{\lambda}^{(0)}$ in Algorithm 3 permits one to use fixed stepsizes and makes the subgradient algorithm [cf.  (17)] evolve along the specified ray yielding constant increments of $g$ as long as $\bm{\lambda}^{(k)}$ resides in an RFGOR region. Intuitively speaking, such a choice is fast in the sense that it directly drives $\bm{\lambda}^{(k)}$ toward $\bm{\lambda}^{\star}$ instead of potential detours where diminishing stepsizes such as $\gamma_{0}/k$, $\gamma_{0}/\sqrt{k}$ can be ineffective.
@@ V-B2 Efficient Communication with FGOR
Note that Algorithm 2 requires each agent $i$ to transmit $n$ real numbers, i.e., $\mathbf{y}_{i}^{(k)}\in{\rm I\!R}^{n}$ in step 5. In addition, it also requires CS to broadcast n(m1)n(m-1) real numbers to agents, cf.  step 8. Therefore, if $b$ bits are used to represent each real number, then Algorithm 2 requires mnb+n(m1)b=nb(2m1)mnb+n(m-1)b=nb(2m-1) bits per epoch  [^3] In the case of $64$-bit double precision floating point number format $b=64$.. However, at the corresponding steps of Algorithm 3 [cf.  step 3 and step 5], only one-bit communication is needed and it does not depend on the length $n$ of $\mathbf{y}_{i}^{(k)}$. More specifically, in each epoch, Algorithm 3 requires only $m+1$ bits, i.e., $m$ bits at step 3 and one bit at step 5, as long as $\bm{\lambda}^{(k)}$ resides in an RFGOR region. It is worth noting that engineering problems in real-world applications can consist of decision variables whose size $n$ can be in the order of several thousands, or even more [Tibshirani_2009]. Thus, by leveraging FGOR characteristics, Algorithm 3 can yield a considerable reduction in the communication overhead.